# Enhancing Compositional Text-to-Image Generation with Reliable Random Seeds

### v3

@@ Abstract
Text-to-image diffusion models have demonstrated remarkable capability in generating realistic images from arbitrary text prompts. However, they often produce inconsistent results for compositional prompts such as “two dogs” or “a penguin on the right of a bowl”. Understanding these inconsistencies is crucial for reliable image generation. In this paper, we highlight the significant role of initial noise in these inconsistencies, where certain noise patterns are more reliable for compositional prompts than others. Our analyses reveal that different initial random seeds tend to guide the model to place objects in distinct image areas, potentially adhering to specific patterns of camera angles and image composition associated with the seed. To improve the model’s compositional ability, we propose a method for mining these reliable cases, resulting in a curated training set of generated images without requiring any manual annotation. By fine-tuning text-to-image models on these generated images, we significantly enhance their compositional capabilities. For numerical composition, we observe relative increases of 29.3% and 19.5% for Stable Diffusion and PixArt-$\alpha$, respectively. Spatial composition sees even larger gains, with 60.7% for Stable Diffusion and 21.1% for PixArt-$\alpha$.
@@ 1 Introduction
Text-to-image synthesis transforms textual descriptions into images, supporting various applications like digital arts, visualization, design, and education. Diffusion models, such as Stable Diffusion (Rombach et al., 2022) and PixArt-$\alpha$ (Chen et al., 2023a), have become the gold standard for generating high-quality, diverse images, far surpassing earlier GAN-based methods (Reed et al., 2016; Liang et al., 2020; Tao et al., 2022). However, these models are not without limitations: They often struggle with compositional prompts, frequently generating inaccurate numerical and spatial details (e.g., ”four unicorns” or ”a dove on top of a basketball”), as shown in Figure 1. To remedy this issue, some works improve the models by incorporating layout inputs for object positioning (Dahary et al., 2024; Chen et al., 2023b; Zheng et al., 2023), while others leverage Large Language Models (LLMs) to generate bounding box coordinates from prompts (Qu et al., 2023; Lian et al., 2023; Feng et al., 2024a; Chen et al., 2023c). These approaches mitigate the issue to a certain extent, but also face various challenges, including labor-intensive layout creation, limited diversity, biases, and potential negative impacts on image realism and generation speed.
Furthermore, the exact reason for the model’s inconsistencies for compositional prompts remain unclear, as most studies focus on understanding the overall performance of the model without addressing specific prompt types. For instance, Guo et al. (2024a) demonstrate that the initial noise significantly influences the quality of the generated images, with certain noise patterns leading to semantically coherent outputs while others not. Similarly, Xu et al. (2024) point out that latent noise vectors located in high-density regions are more likely to yield higher quality samples. These findings commonly suggest a significant role of noise in understanding the underlying mechanisms of diffusion models.
@@ 3.1 Initial Seeds Affect Object Arrangements
We find a strong correlation between initial seeds and object arrangement. Note that object arrangement refers to patterns of object placement in terms of relative positions rather than explicit layouts, such as bounding boxes or region masks, which specify the positions, sizes, and shapes of objects in the image (Chen et al., 2023b; Qu et al., 2023; Zheng et al., 2023; Couairon et al., 2023; Jia et al., 2024; Lian et al., 2023). To evidence this, we used Stable Diffusion 2.1 to generate 512 images based on 8 seeds, 8 classes, and 8 backgrounds, using prompts in the format “Four {object category}, {background setting}” (see Appendix A.1.4). Figure 3 shows the average cross-attention map corresponding to the object token for each seed. As can be seen, each initial seed leads to distinct areas where objects are more likely to be placed.
We provide visual examples in Figure 3 for two seeds: 50 and 23. For seed $50$, it can be seen that the objects tend to be placed in the center diagonally side by side, providing generally more space for the objects. In contrast, for seed $23$, objects often occupy very small areas (e.g., the first two images in the first row) and frequently fail to properly appear in the image. Albeit not perfect, using seed 50 is generally easier than using seed 23 for Stable Diffusion to generate accurate images for compositional prompts. In fact, for seed $50$, 14 out of 16 images are correct, whereas for seed $23$, only 3 images are correct. These distinct patterns highlight the significant impact of initial seeds on the overall object arrangement.
@@ 4 Mining Reliable Seeds for Enhancing Generation
Our analyses reveal that some initial seeds lead to much better compositional performance. To enhance text-to-image models, a straightforward approach is to use only these seeds for image generation. We go one step further and propose to construct a dataset generated from those reliable seeds and use it to improve the overall model performance via fine-tuning.
To this end, we first construct the Comp90 dataset, a prompt dataset for text-to-image generation and reliable seed mining that contains 3,000 text prompts for numerical composition and 3,200 for spatial composition, composed of 90 popular object categories and 12 background settings. We employ the off-the-shelf vision language model CogVLM2 (Hong et al., 2024) for finding reliable seeds. The overview of the proposed approach is illustrated in Figure 5.

### v5

@@ Abstract
Text-to-image diffusion models have demonstrated remarkable capability in generating realistic images from arbitrary text prompts. However, they often produce inconsistent results for compositional prompts such as “two dogs” or “a penguin on the right of a bowl”. Understanding these inconsistencies is crucial for reliable image generation. In this paper, we highlight the significant role of initial noise in these inconsistencies, where certain noise patterns are more reliable for compositional prompts than others. Our analyses reveal that different initial random seeds tend to guide the model to place objects in distinct image areas, potentially adhering to specific patterns of camera angles and image composition associated with the seed. To improve the model’s compositional ability, we propose a method for mining these reliable cases, resulting in a curated training set of generated images without requiring any manual annotation. By fine-tuning text-to-image models on these generated images, we significantly enhance their compositional capabilities. For numerical composition, we observe relative increases of 29.3% and 19.5% for Stable Diffusion and PixArt-$\alpha$, respectively. Spatial composition sees even larger gains, with 60.7% for Stable Diffusion and 21.1% for PixArt-$\alpha$. Our code is available at https://github.com/doub7e/Reliable-Random-Seeds.
@@ 1 Introduction
Text-to-image synthesis transforms textual descriptions into images, supporting various applications like digital arts, visualization, design, and education. Diffusion models, such as Stable Diffusion (Rombach et al., 2022) and PixArt-$\alpha$ (Chen et al., 2023a), have become the gold standard for generating high-quality, diverse images, far surpassing earlier GAN-based methods (Reed et al., 2016; Liang et al., 2020; Tao et al., 2022). However, these models are not without limitations: They often struggle with compositional prompts, frequently generating inaccurate numerical and spatial details (e.g., “four unicorns” or “a dove on top of a basketball”), as shown in Figure 1. To remedy this issue, some works improve the models by incorporating layout inputs for object positioning (Dahary et al., 2024; Chen et al., 2023b; Zheng et al., 2023), while others leverage Large Language Models (LLMs) to generate bounding box coordinates from prompts (Qu et al., 2023; Lian et al., 2023; Feng et al., 2024a; Chen et al., 2023c). These approaches mitigate the issue to some extent, but also face various challenges, including labor-intensive layout creation, limited diversity, biases, and potential trade-offs in image realism and generation speed.
Furthermore, the exact reasons for the model’s inconsistencies with compositional prompts remain unclear, as most studies focus on understanding the overall performance of the model without addressing specific prompt types. For instance, Guo et al. (2024a) demonstrate that the initial noise significantly influences the quality of generated images, with certain noise patterns leading to semantically coherent outputs while others not. Similarly, Xu et al. (2024) highlight that latent noise vectors located in high-density regions are more likely to yield higher-quality samples. These findings commonly suggest a significant role of noise in understanding the underlying mechanisms of diffusion models.
@@ 3.1 Initial Seeds Affect Object Arrangements
We find a strong correlation between initial seeds and object arrangement. Note that object arrangement refers to patterns of object placement in terms of relative positions rather than explicit layouts, such as bounding boxes or region masks, which specify the positions, sizes, and shapes of objects in the image (Chen et al., 2023b; Qu et al., 2023; Zheng et al., 2023; Couairon et al., 2023; Jia et al., 2024; Lian et al., 2023; Xu et al., 2023). To evidence this, we used Stable Diffusion 2.1 to generate 512 images based on 8 seeds, 8 classes, and 8 backgrounds, using prompts in the format “Four {object category}, {background setting}” (see Appendix A.1.4). Figure 3 shows the average cross-attention map corresponding to the object token for each seed. As can be seen, each initial seed leads to distinct areas where objects are more likely to be placed.
We provide visual examples in Figure 3 for two seeds: 50 and 23. For seed $50$, it can be seen that the objects tend to be placed in the center diagonally side by side, providing generally more space for the objects. In contrast, for seed $23$, objects often occupy very small areas (e.g., the first two images in the first row) and frequently fail to properly appear in the image. Although not perfect, using seed 50 is generally more effective than using seed 23 in guiding Stable Diffusion to generate accurate images for compositional prompts. In fact, for seed $50$, 14 out of 16 images are correct, whereas for seed $23$, only 3 images are correct. These distinct patterns highlight the significant impact of initial seeds on the overall object arrangement.
@@ 4 Mining Reliable Seeds for Enhancing Generation
Our analyses reveal that some initial seeds lead to significantly better compositional performance. To enhance text-to-image models, a straightforward approach is to use only these seeds for image generation. We go one step further and propose to construct a dataset generated from those reliable seeds and use it to improve the overall model performance via fine-tuning.
To this end, we first construct the Comp90 dataset– a prompt dataset for text-to-image generation and reliable seed mining– containing 3,000 text prompts for numerical composition and 3,200 for spatial composition, spanning 90 popular object categories and 12 background settings. We employ the off-the-shelf vision language model CogVLM2 (Hong et al., 2024) for finding reliable seeds. The overview of the proposed approach is illustrated in Figure 5.
@@ 6 Conclusion
7 Acknowledgement
This work was partially funded by the Swiss National Science Foundation via Project 200020_214878.